Suppose a security pays $100 in 3 months, and is currently trading at $96.55. What is the effective annual rate on this security?

Round your answer to 4 decimal places. For example if your answer is 3.205%, then please write down 0.0321.
Answer #1

Return = [FV - PV] / PV = [$100 - $96.55] / $96.55 = $3.45 / $96.55 = 0.0357

EAR = [1 + [APR / n]]T x n - 1 = [1 + [0.0357 / 4]]4 - 1 = 1.0362 - 1 = 0.0362